OpenStudy (anonymous):

tell whether each percent change is an increase or decrease. Then find the percent change. Round to the nearest percent. original amount: 45 new amount: 60

OpenStudy (aravindg):

What do you think about this question? Any ideas?

OpenStudy (anonymous):

None. My teacher hasn't taught me this yet. I know it's an increase though.

hero (hero):

\[\%\Delta = \frac{|a - b|}{a} \times 100\] \(\%\Delta = \) Percent Change \(a = \) Old Amount \(b = \) New amount

OpenStudy (anonymous):

Ook Can u use the actual numbers in that ?

hero (hero):

Yes, that's what you're supposed to do. In this case: \(a = 45\) \(b = 60\)

OpenStudy (anonymous):

Thank u. It's hard for me to understand this stuff. ok so it would be x = 45−60 ==== ×100 45 45 - 60 = -15 15 == = 3 45 3(100) = 300

hero (hero):

\[\frac{15}{45} = \frac{1}{3}\]

OpenStudy (anonymous):

ok so 1/3(100) = 33

hero (hero):

Correct.

OpenStudy (anonymous):

Thank u !